OpenStudy (anonymous):

Mr Brown asks a garden designer to make a square flower bed. he cost of paving the border of the flower bed is $20 per metre. The cost of the flowers and soil in the flower bed is $60 per m^2. If the total cost of making the flower bed is $3500, find the length of a side of the flower bed,

OpenStudy (id21):

So, lets tell that one side of this square flowerbed is x. Perimeter of this area is 4x Area is:\[Area=x^2\] Paving the border costs:\[4x \times 20\] The flowers and soil costs: \[x^2 \times 60\] So the equation is: \[60x^2+80x=3500\]

OpenStudy (id21):

\[x_{1,2}=7;-50/6\] Only positive answer fits to you. So the length of this flowerbeds side is 7m

OpenStudy (anonymous):

when do we use quadratic to solve and when do we use simultaneous(elimination and substitution method)?

OpenStudy (id21):

Well, you need to sum both costs to get total, that's how it became quadratic equation.